Ms kanjee who works as a salesperson at a certain company earns 4 1/2 % commission the first R15000 of sales and 7% on any further sales made over and above R200000 during this month,Ms kanjee plans to sell goods that worth R500000 calculate the commission she will receive if she will be able to sell goods as planned
1
Expert's answer
2021-10-21T04:32:00-0400

For the first sale commission earned "=4.5/100\times 15000=R675"

Sales above R200000=500000-200000=R300000

"commission" ="7/100\times 300000=R21000"

total commission =21000+675=R21675